The question say "Angle A and angle B are complementary. If angle A= 3x + 40, and angle B= 2x-10, find the measure of each angle
Answer by stanbon(75887)

The question say "Angle A and angle B are complementary. If angle A= 3x + 40, and angle B= 2x-10, find the measure of each angle
-------
A + B = 90 degrees
-------
3x+40 + 2x-10 = 90
5x = 60
x = 12
------
A = 3x+40 = 3*12+40 = 76 degrees
----
B = 2x-10 = 2*12-10 = = 14 degrees
====================
Cheers,
Stan H.
===============